OpenStudy (anonymous):

Identify the following geometric objects in 3D space: 1) r=2cos (theta) + 2sin (theta) 2) r= 1/(1-cos(theta))

OpenStudy (anonymous):

Convert them: Multiply by r. r^2=2rcos(theta)+2rsin(theta) x^2+y^2=2x+2y So its a circle (complete the square) in a 2-d plane. So its a circular cylinder in 3space.

OpenStudy (anonymous):

Now you do the second one.

OpenStudy (anonymous):

is it r^2 = r( 1/(1-cos(theta))

OpenStudy (anonymous):

and then... x^2 + y^2 = xy/ (1-x)

OpenStudy (anonymous):

i cant quite get my mind around it!

OpenStudy (anonymous):

could you guide me through it please?!

OpenStudy (anonymous):

I'm honestly lost of how to convert it myself. H/o.

OpenStudy (anonymous):

I can't get rid of that r. Its a bit ridiculous. There are some functions that you can't convert. This is the best I can do you man. http://www.wolframalpha.com/input/?i=polarplot+r%3D1%2F%281-cos%28theta%29%29%2Ctheta%2C0%2C8pi This would give you a parabolic"ish" cylinder in 3 space.

OpenStudy (anonymous):

Thanks man.. i appreciate it... here is another one i am confused on... can you take a look? please... http://openstudy.com/groups/mathematics#/users/benfraser1012